1. A mole of sugar cubes stacked on top of one another would reach from here to the sun and back how many times? hsugar cube  = 1.0 cm
  2. The average mass of a mole is 49 g. The moon would have a mass how many times greater than a mole of moles?




1
Expert's answer
2021-03-08T07:07:57-0500

1. The volume of one unit is 1 cm3=1×10−6m3, so the volume of a mole of them is 6.02×1023cm3=6.02×1017m3. The cube root of this number gives the edge length 8.4×105 m3. This is equivalent to roughly 8×102 km.

2. Mass of moon 7.36 × 1022 kg.

7.36 × 1022/0.049 = 1,49591837e24.